OpenStudy (anonymous):

How do I set up this problem. Converting the angle 35' 58' 33' to decimal degrees

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

get a calculator then press this sign\[,,,,\]

OpenStudy (imstuck):

or you could do it the way your teacher wants you to do it by converting using the conversions from minutes and seconds to degrees.

OpenStudy (phi):

do you mean 35º 58' 33" ?

OpenStudy (phi):

there are 60 seconds in one minute change 33" (33 seconds) to a fraction of a minute can you do that ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

\[\frac{ 33 }{ 60 }\] ?

OpenStudy (phi):

here is an example: 30 seconds is 30/60 = ½ of a minute yes 33/60 now change to a decimal.

OpenStudy (anonymous):

.55

OpenStudy (phi):

it's easy with a calculator also, if you practice with numbers, you would know both the top and bottom can be divided by 3, to get 11/20 and if you multiply top and bottom by 5 you get 55/100 and that is easy to change to a decimal. (divide by 100 means move the decimal point t places to the left)

OpenStudy (anonymous):

I'm a bit lost now

OpenStudy (phi):

we know 38" = 0.55' (33 seconds is the same as 0.55 minutes) 35º 58' 33" is the same as 35º 58' + 0.55' or 35º 58.55' now change 58.55' to a fraction of a degree (divide by 60)

OpenStudy (phi):

change minutes to degrees by dividing by 60 (there are 60 minutes in 1 degree) if you had 30 minutes, then you have 30/60 degrees = ½ degree = 0.5 degree if you have 58.55 minutes, then you have 58.55/60 degrees

OpenStudy (anonymous):

I get .96 is that right ?

OpenStudy (phi):

it's (almost) the same thing as working with time. 30 minutes is ½ of an hour 58.55 minutes is practically a whole hour. It would be 58.55/60 ths of an hour 58.55/60= 0.9758333... (the 3 repeats) unless they say how accurate the answer should be, put all of the decimals in now you have 35º + 0.9758333º = 35.97583...º

OpenStudy (phi):

the idea is to do it step by step: change the seconds to minutes and add that fraction to the minutes then change *those* minutes to degrees, and add that to the whole degrees you get degrees as a decimal number.

OpenStudy (phi):

*** I get .96 is that right ? *** that is not quite correct. How did you get it ?

OpenStudy (anonymous):

@phi thank you :)